# Benchmarking Vision Language Model Unlearning via Fictitious Facial Identity Dataset

### v1

@@ 2.5.1 Model Utility
GPT-Eval (GPT). Traditional metrics may overlook semantic information (Wang et al., 2023a), which is crucial for evaluating VLMs. Inspired by LLM-as-a-Judge (Zheng et al., 2024), we introduce GPT-eval, applying GPT-4o mini for model performance evaluation focusing on semantic meanings, which has been commonly used as a metric in evaluating VLMs, such as LLaVA (Liu et al., 2024a) and Video-ChatGPT (Maaz et al., 2023). With the judgment role of GPT-4o mini, GPT-eval focuses on correctness, helpfulness, and relevance, assigning an overall score on a scale of 0 to 1. Additionally, it is tasked with generating a comprehensive explanation of the evaluation, facilitating a better understanding of the predictions. Detailed prompt used in LLM-as-a-Judge for GPT-Eval presenters in Appendix H.
@@ 3.2 Stage II: VLM Unlearning
Unlearning Method Model Utility Forget Quality Rouge↑ GPT↑ Truth.↑ ACC↑ Avg.↑ KS-Test↑ EM↓ MIA↓ APE↓ Avg.↑ LLaVA-Phi-mini-3B Retain Model 93.7 83.3 77.3 100.0 88.6 0.00 1.10 12.3 1.60 97.1 GA (Yao et al., 2023) 50.6 10.0 48.4 60.8 42.5 -0.03 10.5 13.0 15.3 93.5 GD (Liu et al., 2022) 70.2 37.1 59.6 79.5 61.6 -5.80 12.5 12.6 14.7 82.3 KL (Yao et al., 2024a) 90.3 56.7 72.6 93.3 78.2 -21.7 32.3 42.1 32.7 40.3 PO (Maini et al., 2024) 72.8 33.6 60.7 90.6 64.5 -4.70 5.38 64.2 7.62 76.1 LLama-3.2-Vision-Instruct-11B Retain Model 88.8 80.2 72.3 100.0 85.3 0.00 0.81 12.2 1.32 96.1 GA (Yao et al., 2023) 4.30 0.50 42.1 0.00 11.7 -0.85 1.61 15.3 2.47 93.1 GD (Liu et al., 2022) 64.2 23.7 53.8 52.1 48.4 -1.35 16.9 18.2 20.7 87.7 KL (Yao et al., 2024a) 55.6 27.9 65.8 60.5 52.4 -3.75 4.27 13.7 5.35 86.3 PO (Maini et al., 2024) 68.7 42.5 60.0 91.5 65.7 -12.0 1.10 42.3 3.55 59.2
@@ 3.3 Ablation Studies
Fine-tuning VLMs with parameters from different components. Unlike LLMs, VLMs typically consist of three components: a vision encoder, a projector, and an autoregressive LLM. Referring to previous work (Liu et al., 2023; Ye et al., 2023; Dai et al., 2024), we summarize four common fine-tuning strategies for VLMs (from $Ex_{1}$ to $Ex_{4}$), as shown in Table 3. When performing unlearning by fine-tuning various components of VLMs, the results reveal that fine-tuning only the projector would neither achieve the desired forget quality nor maintain good model utility. Fine-tuning both the vision encoder and the projector may change the extracted image features, which facilitates forgetting but significantly impacts model performance on general images. On the other hand, fine-tuning either the projector and LLM, or just the LLM, achieves a more balanced unlearning efficacy. Consequently, for VLM unlearning, we finally decided to fine-tune the parameters from LLM due to its simplicity and effectiveness.
@@ C.1 Kmeans for flitering similar images
Each class in the ISEKAI dataset contains 50-150 images of fictitious entities that were generated by MidJourney Midjourney (2023). this would result in many highly similar images, which is not conducive to scaling the diversity of our benchmark, we first used Kmeans for filtering similar images. Specifically, let $I$ be the set of images in each category of ISEKAI and $f(I)$ be the function that converts an image to its vector representation using CLIP Radford et al. (2021). We employ UMAP McInnes et al. (2018) to further reduce the dimensionality of the CLIP features, followed by a K-means clustering Hartigan & Wong (1979) process with cluster number $k$:

### v2

@@ 2.5.1 Model Utility
GPT-Eval (GPT). Traditional metrics may overlook semantic information (Wang et al., 2023a), which is crucial for evaluating VLMs. Inspired by LLM-as-a-Judge (Zheng et al., 2024), we introduce GPT-eval, applying GPT-4o mini for model performance evaluation focusing on semantic meanings, which has been commonly used as a metric in evaluating VLMs, such as LLaVA (Liu et al., 2024a) and Video-ChatGPT (Maaz et al., 2023). With the judgment role of GPT-4o mini, GPT-eval focuses on correctness, helpfulness, and relevance, assigning an overall score on a scale of 0 to 1. In practice, we multiply the GPT scores by 100 for calculating the reasonable average with the other metrics of model utility. Additionally, it is tasked with generating a comprehensive explanation of the evaluation, facilitating a better understanding of the predictions. Detailed prompt used in LLM-as-a-Judge for GPT-Eval presenters in Appendix H.
@@ 3.2 Stage II: VLM Unlearning
Unlearning Method Model Utility Forget Quality Rouge↑ GPT↑ Truth.↑ ACC↑ Avg.↑ KS-Test↑ EM↓ MIA↓ APE↓ Avg.↑ LLaVA-Phi-mini-3B Retain Model 93.7 83.3 77.3 100.0 88.6 0.00 13.3 12.3 14.7 93.3 GA (Yao et al., 2023) 50.6 10.0 48.4 60.8 42.5 -0.03 14.0 13.0 15.4 92.9 GD (Liu et al., 2022) 70.2 37.1 59.6 79.5 61.6 -5.80 18.9 12.6 19.0 80.7 KL (Yao et al., 2024a) 90.3 56.7 72.6 93.3 78.2 -21.7 41.2 42.1 46.8 36.8 PO (Maini et al., 2024) 72.8 33.6 60.7 90.6 64.5 -4.70 6.90 64.2 6.80 76.0 LLama-3.2-Vision-Instruct-11B Retain Model 88.8 80.2 72.3 100.0 85.3 0.00 14.8 12.2 14.3 93.4 GA (Yao et al., 2023) 4.30 0.50 42.1 0.00 11.7 -0.85 1.60 15.3 1.30 93.2 GD (Liu et al., 2022) 64.2 23.7 53.8 52.1 48.4 -1.40 21.9 18.2 23.4 86.9 KL (Yao et al., 2024a) 55.6 27.9 65.8 60.5 52.4 -3.70 4.50 13.7 11.4 85.7 PO (Maini et al., 2024) 68.7 42.5 60.0 91.5 65.7 -12.0 0.30 42.3 0.50 59.6
@@ 3.3 Ablation Studies
Fine-tuning VLMs with parameters from different components. Unlike LLMs, VLMs typically consist of three components: a vision encoder, a projector, and an autoregressive LLM. Referring to previous work (Liu et al., 2023; Ye et al., 2023; Dai et al., 2024), we summarize four common fine-tuning strategies for VLMs (from $Ex_{1}$ to $Ex_{4}$), as shown in Table 3. When performing unlearning by fine-tuning various components of VLMs, the results reveal that fine-tuning only the LLM can maintain good model utility but fails to achieve the desired forgetting effect. Fine-tuning both the vision encoder and the projector may change the extracted image features, which facilitates forgetting but significantly impacts model performance on general images. On the other hand, fine-tuning either the projector and LLM, or just the projector, achieves a more balanced unlearning efficacy. Consequently, to align with the most commonly used visual instruction tuning process, we finally decided to fine-tune the parameters from the projector and LLM due to its effectiveness.
@@ C.1 Kmeans for flitering similar images
The images are sampled from Part 4 of the SFHQ dataset, which consists of 125,754 high-quality 1024x1024 curated face images. These images were generated using ”inspiration” images sampled from the Stable Diffusion v2.1 Rombach et al. (2021) text-to-image generator with various face portrait prompts. However, since real-world faces are highly diverse, encompassing various elements such as age, gender, hairstyles, and more, randomly selecting images makes it challenging to ensure that the final set of 400 facial images is sufficiently diverse. Therefore, we first used Kmeans for filtering similar images. Specifically, let $I$ be the set of images in part 4 of the SFHQ dataset and $f(I)$ be the function that converts an image to its vector representation using CLIP Radford et al. (2021). We employ UMAP McInnes et al. (2018) to further reduce the dimensionality of the CLIP features, followed by a K-means clustering Hartigan & Wong (1979) process with cluster number $k$:
@@ D.3 Forget Qualtiy
D.4 Model Utility-Forget Quality Curve
Figure 4: Model Utility and Forget Quality trade-off curve on LLaVA-Phi-3-mini. Here we convert the Model Utility, EM, and MIA scores to their difference from 100, and then normalize all their values to between 0 and 1. The higher the Forget quality value, the higher the level of forgetfulness of the unlearned models.
The Figure 4 illustrates the relationship between model utility and forget quality, demonstrating that unlearning algorithms involve a trade-off between these two factors. In Figure 4 (a), we use the exact match (EM) scores to represent forget quality. This metric considers unlearning successful as long as the unlearned model’s responses do not contain private keywords. In contrast, Figure 4 (b) employs membership inference attacks to assess the knowledge retained within unlearned models, thereby determining whether private knowledge has been completely removed. The results reveal that the GD method is the most effective unlearning algorithm when maintaining 60% of model performance. Meanwhile, the KL method achieves the highest level of forgetting when retaining 80% of model performance. However, this method only manages to forget 60% of face-related private knowledge, highlighting the urgent need for more effective unlearning algorithms tailored to vision-language models (VLMs).